# Computationally-Efficient Linear Periodically Time-Variant Digital PLL Modeling Using Conversion Matrices and Uncorrelated Upsampling

### v2

@@ Abstract
This paper introduces a conversion matrix method for linear periodically time-variant (LPTV) digital phase-locked loop (DPLL) phase noise modeling that offers precise and computationally efficient results to enable rapid design iteration and optimization. Unlike many previous studies, which either assume linear time-invariance (LTI) and therefore overlook phase noise aliasing effects, or solve LPTV systems with noise folding and multiple sampling rate conversions that heightens modeling and computational complexity, the proposed conversion matrix method allows the designer to represent the LPTV systems using intuitive LTI-like transfer functions with excellent accuracy. Additionally, computational efficiency is improved through the uncorrelated upsampling method, which eliminates the need to consider beat frequency of noise sources with different sampling rates. The proposed algorithm is applied to modeling a DPLL with time-varying proportional loop gain, and the modeling accuracy is validated with Simulink transient simulations.
@@ I Introduction
Contemporary System-on-Chip (SoC) designs typically require the integration of multiple frequency synthesizers to deliver precise clocks for mixers and digital circuits. Such circuits can end up occupying large physical area (e.g., 25% of the total chip area in [1]), and can end up limiting the performance of such systems. Digital phase-locked loops (DPLLs) are a compelling option for frequency synthesis, due to their technology-scalable area, high programmability, and excellent noise-power figure-of-merit (FoM) [2, 3, 4, 5, 6, 7, 8, 9].
@@
Designing high-performance DPLLs necessitates a comprehensive understanding of the output phase noise (PN). To determine the transfer function from each noise source to the output, many prior studies have treated PLLs as linear time-invariant (LTI) systems at the reference frequency, $f_{ref}$, neglecting noise aliasing caused by divider downsampling [10, 11, 12, 13, 14, 15, 16]. This LTI approximation is suitable for PLLs with relatively low bandwidth. However, for PLLs with very wide bandwidth to suppress the excessive PN from ring oscillators, e.g., 80 MHz in [17] and $\sim 100$ MHz in [18], the aliased noise will contribute significantly to the in-band PN and thus cannot be ignored. Additionally, [12] uses forward Euler mapping $z=1+sT_{ref}$ for $z$-domain to $s$-domain conversion, making the model less accurate in predicting PN close to $f_{ref}/2$.
@@
For a precise modeling of PN in wide-bandwidth DPLLs and multiplying delay-locked loops (MDLLs), it’s essential to consider the system as a discrete-time linear periodically time-variant (LPTV) system. The methodologies in [19, 20, 21] provides great accuracy on their respective models under analysis. However, these math-intensive, topology-specific derivations hinder engineers from quickly adapting to increasingly complicated new topologies [20, 21]. The complexity increases further when a delta-sigma modulator (DSM) with a distinct clock frequency from the oscillator and reference is incorporated [19], necessitating multiple sample-rate conversions within the loop.
@@
The conversion matrix method, also known as the harmonic transfer matrix (HTM) method [22], offers an alternative approach for solving transfer functions in LPTV systems. This approach, previously employed in mixer/$N$-path filter analysis [23, 24, 25], has demonstrated simplified yet accurate results compared to traditional state-space equation solutions [26, 27, 28, 29]. Moreover, it has been leveraged in stability analysis of grid-connected PLLs in power converters [30, 31, 32]. The primary advantage offered by the conversion matrix method is its capability to represent all LPTV components with matrices, where the elements denote the ratio of the input and output frequency-shifted spectrum. Consequently, formulating the problem is straightforward for designers, and computationally demanding matrix operations are offloaded to modern computers optimized for such calculations. The conversion matrix method provides a modeling framework that can be easily applied to different designs with arbitrary complexity.
@@
PLLs, which are inherently LPTV, appear to be an ideal candidate for the conversion matrix method. However, the incorporation of a DSM could significantly increase the computational complexity. To understand why, consider simulating the conversion gain of an $RC$-mixer using Spectre periodic steady-state (PSS) and periodic AC (PAC) analysis. When the clock frequency, $f_{s}$, is 1 GHz, the fundamental frequency in PSS equals to $f_{s}=1$ GHz; the simulation can quickly converge and only a few sidebands, depending on the application, need to be saved. However, incorporating an additional mixer with a clock frequency $f_{s}^{{}^{\prime}}=f_{s}+\Delta f=$ 1.01 GHz changes the system’s fundamental frequency to the beat frequency $\Delta f=$ 0.01 GHz. As a result, 100$\times$ more sidebands need to be saved, significantly increasing the simulation memory and time consumption. Similar challenges arise when using the conversion matrix method to model a PLL if the DSM frequency is not an integer multiple of the reference frequency. Here, the beat frequency becomes the fundamental frequency, potentially making the problem dimension of the conversion matrix method (similar to the number of sidebands in PSS) too large to be solved.
@@
    Introducing uncorrelated upsampling that allows ignoring any beat frequency between different-rate noise sources, thereby reducing computation time.
@@
The paper is organized as follows: Section II defines the conversion matrix method in discrete-time LPTV systems. Section III introduces uncorrelated upsampling and provides an example of modeling a PLL with time-varying proportional loop gain. Section IV compares the simulation results with the PN predicted by the model. Finally, Section V offers conclusions.
@@ II Discrete-Time LPTV System Modeling
In a discrete-time LPTV system depicted in Fig. 1, the relationship between input $x[n]$ and output $y[n]$ is characterized by the time-varying impulse responses of the feed-forward and feedback paths, represented as $p_{n}[n]$ and $f_{n}[n]$, respectively. The LPTV system’s frequency conversion behavior, with a sampling rate of $f_{s}=1/T_{s}$, can be eventually expressed first by using a vector, $\mathbf{X}$, composed of the frequency-shifted input spectrum:
@@
in which $X(\Omega)$ is the discrete-time Fourier transform (DTFT) of $x[n]$ [34, eq. (12.1)], $\Omega=2\pi fT_{s}$ is the discrete angular frequency, $K$ is defined as the problem dimension, $k$ is the harmonic index of the input, and $2\pi/K$ is the fundamental discrete angular frequency of the time-varying impulse response.
@@
is defined as conversion matrix that relates frequency-shifted input and output spectrum. The elements of $\mathbf{H}$, denoted as $H_{q,k}$, represent the gain from the $k^{\text{th}}$ input harmonic to the $q^{\text{th}}$ output harmonic. The discrete-time conversion matrix has a finite $K\times K$ dimension, as opposed to being a doubly infinite matrix in the continuous-time case [22, eq. (12)]. This is because the DTFT of a discrete-time signal is $2\pi$-periodic, making the definitions in (1)-(4) contain all possible combinations of the input and output harmonics.
@@
$S_{yy}(\Omega)=\sum_{k=0}^{K-1}|H_{0,k}(\Omega)|^{2}Sxx\left(\Omega-\frac{2\pi k}{K}\right).$ (5)
@@
$\mathbf{Y}=[(\mathbf{I}+\mathbf{P}\mathbf{F})^{-1}\mathbf{P}]\mathbf{X}=\underbrace{[(\mathbf{I}+\mathbf{P}\mathbf{F})\backslash\mathbf{P}]}_{\mbox{$\equiv\mathbf{H}$}}\mathbf{X},$ (7)
[…]
@@ II-B Multiplication with Periodic Sequence
Multiplying the input, $x[n]$, with a periodic sequence, $w[n]$, shown in Fig. 3 is another fundamental block in LPTV systems. The sequence $w[n]$ has a period of $N$, i.e., $w[n]=w[n+N]$. Define $w_{0}[n]$ as one period of $w[n]$ [34, eq. (12.17)]
@@
in which $W_{0}(\Omega)$ is the DTFT of $w_{0}[n]$ [34, eq. (12.1)]
@@
Multiplication in the time domain translates to circular convolution in frequency domain [34, eq. (12.14)]. As a result, the output spectrum $Y(\Omega)$ can be written as
@@
Like its continuous-time counterpart, $\mathbf{H}_{\text{Mul}}$ is a $K\times K$ frequency-independent Toeplitz matrix [22, eq. (27)].
@@ III PLL Modeling Example
With the derived building blocks, this section models a digital PLL with fast phase-error correction (FPEC) [7] as an example. In contrast to a conventional PLL that maintains a constant proportional loop gain $K_{P}$, FPEC boosts $K_{P}$ at the beginning of each reference cycle to rapidly eliminate the residue jitter from the previous reference cycle [7, Fig. 2], resulting in reduced average jitter. This is a time-varying concept, and cannot be modeled with conventional LTI techniques.
@@
$K_{P}[n]=\begin{cases}\frac{NK_{P0}}{P},&kN\leq n\leq kN+P-1\\ {0},&{\text{else}}\end{cases},$ (16)
in which $k\in\{0,1,2,\ldots\}$, $P\in[1,N]$ is the length of the FPEC window, and $K_{P0}$ is the proportional loop gain when FPEC is deactivated, i.e., $P=N$. When $P$ decreases, the height of $K_{P}[n]$ increases while keeping the total area within one reference period the same. Subsequent subsections will derive the DCO-rate transfer functions for noise sources with different sampling rates.
@@ III-A Noise Sampling Rate Conversion
    DCO noise, $\phi_{n,DCO}[n]$, running at $f_{DCO}$, which is the highest rate in the model.
@@
    DSM quantization noise, $q_{n,DSM}[p]$, running at $f_{DSM}=f_{DCO}/M$. In most cases $1\leq M<N$ for pushing the shaped quantization noise to higher frequency.
@@
    Reference noise, $\phi_{n,ref}[m]$, running at $f_{ref}=f_{DCO}/N$, which ideally should be the fundamental frequency of the model, making $K=N$ in (1). The phase detector (PD) quantization noise can be lumped into the reference noise by dividing PD gain $K_{PD}$.
@@
The sample-rate upconversion is achieved through uncorrelated upsampling. Uncorrelated upsampling by a factor of $L$ involves upsampling by inserting $L-1$ zeros between adjacent samples of the input $x[m]$, followed by multiplying a decorrelation sequence, $w_{U,L}[n]$, that removes correlated frequency spectrum with respect to the fundamental discrete angular frequency, $2\pi/N$. The PSD of the upsampled sequence $y_{corr}[n]$, $S_{yy,corr}(\Omega)$, can be written as
$S_{yy,corr}(\Omega)=\frac{1}{L}S_{xx}\biggl{(}\text{mod}(L\Omega,2\pi)\biggr{)},\Omega\in[0,2\pi),$ (17)
in which $\text{mod}(a,b)$ finds the remainder after division of $a$ by $b$. One period of $w_{U,L}[n]$ is expressed as
$w_{U0,L}[n]=\begin{cases}\sqrt{\mathrm{gcd}(L,N)},&\text{if mod}(n,\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases},$ (18)
in which $10\leq n\leq N-$ 1, while $\mathrm{gcd}(L,N)$ computes the greatest common divisor of $L$ and $N$. The decorrelation sequence $w_{U,L}[n]$ is the $N$-periodic extension of $w_{U0,L}[n]$:
$w_{U,L}[n]=w_{U0,L}[\text{mod}(n,N)],$ (19)
and its associated conversion matrix $\mathbf{W}_{U,L}$ can be found with (15). An intuitive explanation of $w_{U,L}[n]$ is shown in the Appendix.
Note that multiplying the upsampled noise source with a decorrelation sequence $w_{U,L}[n]$ does not occur in an actual PLL; it only has mathematical significance that allows using $f_{ref}$ as the fundamental frequency of the system. Otherwise, any beat frequency lower than $f_{ref}$ becomes the fundamental frequency, unnecessarily enlarging the problem dimension as $K$ increases in (1). For example, if $N=51$ and $M=50$, the problem dimension $K=\text{lcm}(M,N)=2550$, where lcm represents the least common multiple. Considering the $O(n^{3})$ complexity when inverting an $n\times n$ matrix, the proposed uncorrelated upsampling reduces the computation time by $(\text{lcm}(M,N)/N)^{3}=125000\times$.

### v3

# Linear Periodically Time-Variant Digital PLL Phase Noise Modeling Using Conversion Matrices and Uncorrelated Upsampling

@@ Abstract
This paper introduces a conversion matrix method for linear periodically time-variant (LPTV) digital phase-locked loop (DPLL) phase noise modeling that offers precise and computationally efficient results to enable rapid design iteration and optimization. Unlike many previous studies, which either assume linear time-invariance (LTI) and therefore overlook phase noise aliasing effects, or solve LPTV systems with noise folding and multiple sampling rate conversions that heightens modeling and computational complexity, the proposed conversion matrix method allows the designer to represent LPTV systems using intuitive LTI-like transfer functions with excellent accuracy. Additionally, the uncorrelated upsampling method addresses the cross-correlated spectrum of cyclostationary noise sources by a simple matrix multiplication. This eliminates the need to consider the beat frequency of the upsampled noise source and the system with different sampling rates, thus improving computational efficiency. The proposed algorithm is applied to modeling a integer-$N$ DPLL with time-varying proportional loop gain, and the modeling accuracy is validated with Simulink transient simulations.
@@ I Introduction
Contemporary System-on-Chip (SoC) designs typically require the integration of multiple frequency synthesizers to deliver precise clocks for mixers and digital circuits. Such circuits can end up occupying large physical area (e.g., 25% of the total chip area in [1]), and can limit the performance of such systems. Digital phase-locked loops (DPLLs) are a compelling option for frequency synthesis, due to their technology-scalable area, high programmability, and excellent noise-power figure-of-merit (FoM) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@
Designing high-performance DPLLs necessitates a comprehensive understanding of the output phase noise (PN). To determine the transfer function from each noise source to the output, many prior studies have treated PLLs as linear time-invariant (LTI) systems at the reference frequency, $f_{\text{REF}}$, neglecting noise aliasing caused by divider downsampling [13, 14, 15, 16, 17, 18, 19]. This LTI approximation is suitable for PLLs with relatively low bandwidth. However, for PLLs with very wide bandwidth to suppress the excessive PN from ring oscillators, e.g., 80 MHz in [20] and $\sim 100$ MHz in [21], the aliased noise will contribute significantly to the in-band PN and thus cannot be ignored. Additionally, [15] uses forward Euler mapping $z=1+sT_{\text{REF}}$ for $z$-domain to $s$-domain conversion, making the model less accurate in predicting PN close to $f_{\text{REF}}/2$.
@@
For precise modeling of PN in wide-bandwidth DPLLs and multiplying delay-locked loops (MDLLs), it’s essential to consider the system as a discrete-time linear periodically time-variant (LPTV) system. The methodologies in [22, 23, 24, 25] provides great accuracy on their respective models under analysis. However, these math-intensive, topology-specific derivations hinder engineers from quickly adapting to increasingly complicated new topologies. The complexity increases further when a delta-sigma modulator (DSM) with a distinct clock frequency from the oscillator and reference is incorporated [23], necessitating multiple sample-rate conversions within the loop. More importantly, these works can produce inaccurate results by ignoring the cross-correlation of the frequency-shifted spectrum, as the upsampled noise sequence is cyclostationary rather than wide-sense stationary.
@@
The conversion matrix method, also known as the harmonic transfer matrix (HTM) method [26], offers an alternative approach for solving transfer functions in LPTV systems with spectral correlation included. This approach, previously employed in mixer/$N$-path filter analysis [27, 28, 29], has demonstrated simplified yet accurate results compared to traditional state-space equation solutions [30, 31, 32, 33]. Moreover, it has been leveraged in stability analysis of grid-connected PLLs in power converters [34, 35, 36]. The primary advantage offered by the conversion matrix method is its capability to represent all LPTV components with matrices, where the elements denote the ratio of the input and output frequency-shifted spectrum. Consequently, formulating the problem is straightforward for designers, and computationally demanding matrix operations are offloaded to modern computers optimized for such calculations. The conversion matrix method provides a modeling framework that can be easily applied to different designs with arbitrary complexity.
@@
PLLs, which are inherently LPTV, appear to be an ideal candidate for the conversion matrix method. However, the incorporation of a DSM could significantly increase the computational complexity. To understand why, consider simulating the conversion gain of an $RC$-mixer using Spectre periodic steady-state (PSS) and periodic AC (PAC) analysis. When the clock frequency, $f_{\text{s}}$, is 1 GHz, the fundamental frequency in PSS equals to $f_{\text{s}}=1$ GHz; the simulation can quickly converge and only a few sidebands, depending on the application, need to be saved. However, incorporating an additional mixer with a clock frequency $f_{\text{s}}^{\prime}=f_{\text{s}}+\Delta f=$ 1.01 GHz changes the system’s fundamental frequency to the beat frequency $\Delta f=$ 0.01 GHz. As a result, 100$\times$ more sidebands need to be saved, significantly increasing the simulation memory and time consumption. Similar challenges arise when using the conversion matrix method to model a PLL if the DSM frequency is not an integer multiple of the reference frequency. Here, the beat frequency becomes the fundamental frequency, potentially making the problem dimension of the conversion matrix method (similar to the number of sidebands in PSS) too large to be solved.
@@
    Introducing uncorrelated upsampling that compensates for the cross-correlation of the frequency-shifted upsampled spectrum. It also allows ignoring any beat frequency between different-rate noise sources, thereby reducing computation time.
@@
The paper is organized as follows: Section II defines the conversion matrix method in discrete-time LPTV systems. Section III introduces uncorrelated upsampling and provides an example of modeling a PLL with time-varying proportional loop gain. Section IV compares the simulation results with the PN predicted by the model. Section V discusses the limitations of the proposed work. Finally, Section VI offers conclusions.
@@ II Discrete-Time LPTV System Modeling
In a discrete-time LPTV system depicted in Fig. 1, the relationship between input $x[n]$ and output $y[n]$ is characterized by the time-varying impulse responses of the feed-forward and feedback paths, represented as $p_{n}[n]$ and $f_{n}[n]$, respectively. The LPTV system’s frequency conversion behavior, with a sampling rate of $f_{\text{s}}=1/T_{\text{s}}$, can be expressed by a vector, $\mathbf{X}$, composed of the frequency-shifted input spectrum:
@@
in which $X(\Omega)$ is the discrete-time Fourier transform (DTFT) of $x[n]$ [38, eq. (12.1)], $\Omega=2\pi fT_{\text{s}}$ is the discrete angular frequency, $K$ is defined as the problem dimension, $k$ is the harmonic index of the input, and $2\pi/K$ is the fundamental discrete angular frequency of the time-varying impulse response.
@@
is defined as conversion matrix that relates frequency-shifted input and output spectrum. The elements of $\mathbf{H}$, denoted as $H_{q,k}$, represent the gain from the $k^{\text{th}}$ input harmonic to the $q^{\text{th}}$ output harmonic. The discrete-time conversion matrix has a finite $K\times K$ dimension, as opposed to being a doubly infinite matrix in the continuous-time case [26, eq. (12)]. This is because the DTFT of a discrete-time signal is $2\pi$-periodic, making the definitions in (1)-(4) contain all possible combinations of the input and output harmonics.
@@
$S_{yy}(\Omega)=\sum_{k=0}^{K-1}|H_{0,k}(\Omega)|^{2}S_{xx}\left(\Omega-\frac{2\pi k}{K}\right).$ (5)
@@
$\mathbf{Y}=\underbrace{[(\mathbf{I}+\mathbf{P}\mathbf{F})^{-1}\mathbf{P}]}_{\mbox{$\equiv\mathbf{H}$}}\mathbf{X},$ (7)
in which $\mathbf{I}$ is a $K\times K$ identity matrix and $(\cdot)^{-1}$ denotes matrix inversion. Note that $\mathbf{H}$ in (7) has a similar form as an LTI system transfer function with the loop gain defined as $\mathbf{L}=\mathbf{P}\mathbf{F}$. When implementing the algorithm, the direction of matrix multiplication needs to be consistent, as matrix multiplication is not commutative in most cases, i.e., $\mathbf{P}\mathbf{F}\neq\mathbf{F}\mathbf{P}$. With the definitions in (1)-(4), signals always “come in” from the right side of the conversion matrix and “leave from” the left side.
@@ II-B Multiplication with Periodic Sequence
Multiplying the input, $x[n]$, with a periodic sequence, $w[n]$, shown in Fig. 3 is another fundamental block in LPTV systems. The sequence $w[n]$ has a period of $N$, i.e., $w[n]=w[n+N]$. Define $w_{0}[n]$ as one period of $w[n]$ [38, eq. (12.17)]
@@
in which $W_{0}(\Omega)$ is the DTFT of $w_{0}[n]$ [38, eq. (12.1)]
@@
Multiplication in the time domain translates to circular convolution in frequency domain [38, eq. (12.14)]. As a result, the output spectrum $Y(\Omega)$ can be written as
@@
Like its continuous-time counterpart, $\mathbf{H}_{\text{MUL}}$ is a $K\times K$ frequency-independent Toeplitz matrix [26, eq. (27)].
@@ III PLL Modeling Example
With the derived building blocks, this section models a digital PLL with fast phase-error correction (FPEC) [7] as an example. In contrast to a conventional PLL that maintains a constant proportional loop gain $K_{\text{P0}}$, FPEC boosts $K_{P}$ at the beginning of each reference cycle to rapidly eliminate the residue jitter from the previous reference cycle [7, Fig. 2], resulting in reduced average jitter. This is a time-varying concept, and cannot be modeled with conventional LTI techniques.
@@
$K_{P}[n]=\begin{cases}\frac{NK_{\text{P0}}}{P},&kN\leq n\leq kN+P-1\\ {0},&{\text{else}}\end{cases},$ (16)
in which $k\in\{0,1,2,\ldots\}$, $P\in[1,N]$ is the length of the FPEC window, and $K_{\text{P0}}$ is the proportional loop gain when FPEC is deactivated, i.e., $P=N$. When $P$ decreases, the height of $K_{P}[n]$ increases while keeping the total area within one reference period the same. Subsequent subsections will derive the DCO-rate transfer functions for noise sources with different sampling rates.
The second-order DSM used in this work adopts the structure depicted in [40, Fig. 8] and connects its input to the integral path similar to [7, Fig. 10] to enhance the DCO tuning resolution. The DSM has 2 delays for the input control code and 2 zeros at the origin for the quantization noise. Unless otherwise mentioned, 2 simplifications relative to [7] are employed:
  1. 1.
    The proportional path and integral path directly control the capacitor banks with the same resolution digitally.
  2. 2.
    All bits in the integral path are connected to the DSM input. Fractional resampling and 2 delays in the signal transfer function (STF) are ignored.
The second simplification may lead to inaccurate results because the fractional resampling of the control code [23] and the delays are ignored. However, the introduced error is negligible because $K_{I}\ll K_{P}$ in practical cases. To avoid distracting the reader from the main framework, the second simplification is made and will be revisited in Section IV-E.
@@ III-A Noise Sampling Rate Conversion
    DCO noise, $\phi_{\text{n,DCO}}[n]$, running at $f_{\text{DCO}}$, which is the highest rate in the model.
@@
    DSM quantization noise, $q_{\text{n,DSM}}[p]$, running at $f_{\text{DSM}}=f_{\text{DCO}}/M$. In most cases $1\leq M<N$ for pushing the shaped quantization noise to higher frequency.
@@
    Reference noise, $\phi_{\text{n,REF}}[m]$, running at $f_{\text{REF}}=f_{\text{DCO}}/N$, which ideally should be the fundamental frequency of the model, making $K=N$ in (1). The phase detector (PD) quantization noise can be lumped into the reference noise by dividing PD gain $K_{\text{PD}}$.
@@
Figure 5: Cyclostationary noise introduced by upsampling a WSS noise sequence by $M$ times. Beat frequency can be generated in the $N$-periodic system if $N$ is not an integer multiple of $M$. The zero-order hold is ignored only for simpler illustration of the beat frequency; it’s included in the final model in Fig. 7.
Proper handling of sample-rate conversion is crucial for calculating transfer functions in LPTV systems. An example of upsampling the DSM noise, $q_{\text{n,DSM}}[p]$, is shown in Fig. 5 with $M=3$ and $N=4$. After upsampling, $q_{\text{n,DSM,up,corr}}[n]$ is no longer WSS but $M$-cyclostationary, i.e., its mean $\mu[n]=\mu[n+lM]$ and autocorrelation sequence $R_{qq}[n,k]=R_{qq}[n+lM,k]$ are periodic over $M$, where $l$ is an integer. When $q_{\text{n,DSM,up,corr}}[n]$ passes through an $N$-periodic system, the PSD of the output, $\phi_{\text{out}}[n]$, cannot be calculated by (5) anymore because it only applies to WSS input that doesn’t have cross-correlated frequency-shifted spectrum [26, eq. (23)].
However, previous works on LPTV PLL analysis use (5) on a cyclostationary input while still demonstrating excellent matching between calculations and simulations. Based on observation, (5) generates no or little error for an $N$-periodic system with $M$-cyclostationary input in one of the following cases:
  1. 1.
    $N$ and $M$ are coprime, i.e., $\text{gcd}(N,M)=1$. In the example showed earlier, $N=4$ and $M=3$, (5) can be used because $\text{gcd}(3,4)=1$. In prior works that model DSM noise, [22] uses $N=25$ and $M=4$, and [23] uses $N=43$ which is a prime number. This is similar to the continuous-time case described in [41, III-A].
  2. 2.
    $M=N$, and the $M$-cyclostationary input comes from upsampling a WSS white noise sequence that has a flat time-averaged PSD.
@@
The special case 1) has an interesting interpretation: when $\text{gcd}(N,M)=1$, the $N$-periodic system treats its $M$-cyclostationary input as WSS, so (5) applies. In other words, the system can’t distinguish whether its input is WSS or not. To generalize special case 1) to all $M$ and $N$ combinations, the sample-rate upconversion in this work is achieved through uncorrelated upsampling, as shown in Fig. 6. Uncorrelated upsampling by a factor of $L$ involves upsampling by inserting $L-1$ zeros between adjacent samples of the input $x[m]$, followed by multiplying a decorrelation sequence, $w_{\text{U},L}[n]$, that removes correlated frequency spectrum with respect to the fundamental discrete angular frequency, $2\pi/N$. It “tricks” the $N$-periodic system into interpreting its input as WSS regardless of $L$.
The PSD of the upsampled sequence $y_{\text{corr}}[n]$, $S_{yy,\text{corr}}(\Omega)$, can be written as
$S_{yy,\text{corr}}(\Omega)=\frac{1}{L}S_{xx}\biggl{(}\text{mod}(L\Omega,2\pi)\biggr{)},\Omega\in[0,2\pi),$ (17)
in which $\text{mod}(a,b)$ finds the remainder after division of $a$ by $b$. One period of $w_{\text{U},L}[n]$ is expressed as
$w_{\text{U0},L}[n]=\begin{cases}\sqrt{\mathrm{gcd}(L,N)},&\text{if mod}(n,\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases},$ (18)
in which $10\leq n\leq N-$ 1, and $\mathrm{gcd}(L,N)$ computes the greatest common divisor of $L$ and $N$. The decorrelation sequence $w_{\text{U},L}[n]$ is the $N$-periodic extension of $w_{\text{U0},L}[n]$:
$w_{\text{U},L}[n]=w_{\text{U0},L}[\text{mod}(n,N)],$ (19)
and its associated conversion matrix $\mathbf{W}_{\text{U},L}$ can be found with (15). An intuitive explanation of $w_{\text{U},L}[n]$ is shown in the Appendix. Note that multiplying the upsampled noise source with a decorrelation sequence $w_{\text{U},L}[n]$ does not occur in an actual PLL; it only has mathematical significance, compensating for the underestimated output PSD due to the cross-correlated frequency-shifted spectrum of the cyclostationary input.
Another practical problem that uncorrelated upsampling solves is that it allows using $f_{\text{REF}}$ as the fundamental frequency of the system. This is because the cyclostationary input appears WSS to the system after decorrelation. Otherwise, any beat frequency lower than $f_{\text{REF}}$ becomes the fundamental frequency, as shown in Fig. 5, unnecessarily enlarging the problem dimension as $K$ increases in (1) to capture all possible frequency conversions. For example, if $N=51$ and $M=50$, the problem dimension $K$ becomes $\text{lcm}(M,N)=2550$, where lcm represents the least common multiple. Considering the $O(n^{3})$ complexity when inverting an $n\times n$ matrix, the proposed uncorrelated upsampling reduces the computation time by $(\text{lcm}(M,N)/N)^{3}=125000\times$.